• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name pzhu

    Log entry time 08:12:30 on March31,2012

    Entry number 368555

    keyword=owl shift summary

    0:35 move target to do dilution run
    1:07 dilution run done, move target back to top NH3
    kalyan changed the beam position to check the rate in 3rd arm
    1:34 controlled access for the target vacuum pump check
    2:30 back to beam permit,changed septum current matched to 1.667GeV P0
    2:40 kalyan changed the beam position to check the rate in 3rd arm, 3rd arm rate is higher than before.
    3:15 start production run for 1.667GeV/c setting
    5:47 bleed-through measurement

    production run summary for 1.667GeV/c:
    L 3519 3520 3521 3522 3523 3524 3525 3526
    R 22551 22552 22553 22554 22555 22558 22559 22560